# Using diffusion model as constraint: Empower Image Restoration Network Training with Diffusion Model

### v1

@@ Abstract
Image restoration has made marvelous progress with the advent of deep learning. Previous methods usually rely on designing powerful network architecture to elevate performance, however, the natural visual effect of the restored results is limited by color and texture distortions. Besides the visual perceptual quality, the semantic perception recovery is an important but often overlooked perspective of restored image, which is crucial for the deployment in high-level tasks. In this paper, we propose a new perspective to resort these issues by introducing a naturalness-oriented and semantic-aware optimization mechanism, dubbed DiffLoss. Specifically, inspired by the powerful distribution coverage capability of the diffusion model for natural image generation, we exploit the Markov chain sampling property of diffusion model and project the restored results of existing networks into the sampling space. Besides, we reveal that the bottleneck feature of diffusion models, also dubbed h-space feature, is a natural high-level semantic space. We delve into this property and propose a semantic-aware loss to further unlock its potential of semantic perception recovery, which paves the way to connect image restoration task and downstream high-level recognition task. With these two strategies, the DiffLoss can endow existing restoration methods with both more natural and semantic-aware results. We verify the effectiveness of our method on substantial common image restoration tasks and benchmarks. Code will be available at https://github.com/JosephTiTan/DiffLoss.
@@ 1 Introduction
In complex imaging environments, the quality of imaging often suffers from unpredictable degradations, such as low-light conditions, heterogeneous media, leading to information loss both in content and color and unnatural visual effects. Besides, these degradations also causes adverse effects on high-level computer vision tasks such as object detection [8, 32] and scene understanding [53, 52]. Thus, it is indispensable to develop effective algorithms to recover high-quality degradation-free images in natural appearance, which is referred to as image restoration. This task has been a longstanding and challenging problem, drawing the attention of researchers.
In fact, estimating the clean image from the degraded input is an ill-posed and challenging problem. Previous traditional methods utilized the inherent feature information of degraded images to design algorithms with prior knowledge for image restoration[4, 23, 16, 15]. However, these hand-crafted image priors are drawn from specific observations with limited robustness, which may not be reliable for modeling the intrinsic characteristics of degraded images. With the advent of deep learning, image restoration has made marvelous progress, where most of them rely on designing advanced architectures to learn degradation to clean mappings [31, 38, 14, 64, 21, 67]. However, due to the severe degradation, the naturalness of the restored results is limited with color and texture distortions, as shown in Fig. 1. This is challenging, if not impossible, to solve in the current framework, wherein we employ mapping-based deep learning paradigm and limited area data. Besides, due to the common practice of relying solely on visual quality metrics during the restoration process, the semantic perception recovery is usually absent in the recovered image, resulting in poor performance in downstream high-level tasks even with seemingly high visual quality.
@@
Figure 1: Visual comparison between baselines with and without our DiffLoss. (a) is produced with IAT[11] on LOL dataset[61]; (b) is produced with MSBDN[14] on NH-HAZE dataset [2]. Without our DiffLoss, the visual result suffers from unnaturalness problem, with color shift and content artifacts. DiffLoss largely alleviates this issue with more natural results. Need to mention that DiffLoss is an optimization strategy. The improvement should be compared with the baseline method, instead of other restoration methods or the ground-truth images.
To this end, we argue that, to meet the challenging information loss issue caused by severe degradation and limited data and model capacity, image restoration methods should be explicitly empowered with models that has strong prior on natural and clean information distribution modeling. Inspiringly, the emerged diffusion model has exhibited strong natural image generation capability [13] and possible semantic potential [3], which open the avenue to solve these problems.
Figure 2: Illustration of images after modifying h-space features with SVD. As h-space changes, the semantic information of the image transitions from an axolotl to a cat and gradually loses its original semantics completely. DiffLoss preserves the original semantic information of the image through h-space.
In this paper, we propose a new perspective to resort these issues by introducing a naturalness-oriented and semantic-aware optimization mechanism, dubbed DiffLoss. Specifically, drawing inspiration from the diffusion model’s remarkable capacity to cover distributions in natural image generation, we leverage the Markov chain sampling characteristic of the diffusion model to project the restored results of existing networks into the sampling space. Besides, we reveal that the bottleneck feature of diffusion models, also dubbed h-space feature, is a natural high-level semantic space, as shown in Fig. 2. We investigate this property and introduce a semantic-aware loss to further unleash its potential in recovering semantic perception. This breakthrough paves the way for bridging the gap between image restoration tasks and downstream high-level recognition tasks. With these two strategies, the DiffLoss can endow existing restoration methods with both more natural and semantic-aware results.
@@
Different from previous methods that focus on dedicated model design or directly employ diffusion model as the restoration model, our DiffLoss works as a general and novel auxiliary training mechanism, to boost the ability of existing methods to achieve better visibility effects and preservation of semantic information. Specifically, we employ the unconditional diffusion model pre-trained on ImageNet dataset as optimization prior to exploit its immense clean data distribution. The diffusion model takes the restored image and ground truth as inputs and projects them into the distribution sampling space and semantic h-space. Within these spaces, we can get more natural result and preserve more semantic information by pulling the intermediate sampling distribution and h-space of the restored image closer to that of the corresponding clear image.
Besides, it’s especially beneficial for empowering parameter-limited models as it involves naturalness prior. Our method also involves no additional computations in the inference stage. We verify the effectiveness of our method on substantial common image restoration tasks, including image dehazing, image deraining, and low-light image enhancement.
@@
  • We introduce the naturalness and recognition-aware modeling paradigm and into restoration network to alleviate the unnaturalness problem and recognition-unfriendly issue of existing methods by embedding the diffusion model as auxiliary training mechanism.
  • Specifically, our approach leverages the fixed diffusion model to enable tractable sampling and the extraction of semantic information, and it yields a natural sample and recognition-aware restoration when the DiffLoss is minimized.
  • Extensive experiments demonstrate that our DiffLoss empowers existing restoration methods and helps training efficient models and improving the Classification ability on data with varying of degradation without involving additional computations in the inference stage.
@@ 2 Related work
2.1 Image Restoration
Image restoration refers to the process of improving the quality of a degraded or damaged image by removing various types of degradations. With the advent of deep learning, numerous methods leveraging deep neural networks for image restoration have emerged, such as low-light enhancement[60, 69, 70, 73], super-resolution[30, 37, 71, 72], dehazing[4, 7, 15, 23, 33, 31, 68, 66], deraining[19, 35, 40, 45, 48], deblurring[5, 9, 20, 29, 49]. However, the severe degradation limits the naturalness of restored results, causing color and texture distortions. Moreover, relying solely on visual quality metrics during the restoration process, without considering the semantic aspects of quality, has negatively impacted performance in downstream tasks.
@@ 2.2 Diffusion models for image restoration
However, all these methods attempt to integrate diffusion models into the network architecture that has the limit of slow inference speed and accounting huge memory cost when deploying. We circumvent these limitations from a new perspective by exploring diffusion model as an auxiliary training mechanism to empower the learning capability of existing image restoration networks, without involving additional computations in the inference stage.
3 Motivation
Inspiration from 3D generation. Our DiffLoss is similar to the Score Distillation Sampling (SDS) in Dreamfusion  [44] that perform text-to-3D synthesis with NeRF model. By using a pretrained 2D text-to-image diffusion model to produce a prior optimization loss and generate corresponding 3D object given a text, the NeRF model distills prior information of natural images from the diffusion model. Similarly, we leverage the unconditional diffusion model pretrained on large-scale ImageNet dataset as optimization prior. Different from SDS, our method is exploited in the 2D image restoration tasks instead of 3D generation and our method is a supervised regression task with degraded image as input and clear image as supervision, while Dreamfusion is a zero-shot generation task with only text prompt as input.
Why DiffLoss can empower restoration networks. Previous loss [12, 17] that assist restoration networks have the following drawbacks: (1) L1 or L2 loss works in pixel space, which may produce images deviated from natural distribution. (2) The VGG16 used in perceptual loss is pre-trained for high-level task, instead of low-level image restoration task. (3) Adversarial loss treats restoration network as generator and inserts an additional discriminator network and needs to train a discriminator for every restoration dataset, which is troublesome and time-consuming.
Thanks to the strong generation ability of DDPMs and the immense data samples in the training dataset, the unconditional diffusion model pre-trained on ImageNet dataset possesses world-wide natural image data distribution and has been applied in image restoration tasks  [39, 27, 58, 55, 41]. The wide distribution and mode convergence property enables it to be a powerful and general image prior, and fits for both general and specific low-level image restoration tasks. Besides, h-space in diffusion model is also found to reflect the semantics of images. Hence, by pulling the intermediate sampling stages and h-space closer to that of clear images and leveraging the distribution sampling property of diffusion model, the restored results can be optimized to be more natural and recognition-aware, which is difficult to achieve for the previous methods. In contrast to previous usage of unconditional diffusion model in low-level vision usually directly exploits its generation ability by incorporating diffusion model into the network backbone, our method propose the "implicit" usage of unconditional diffusion model and employ it as auxiliary training mechanism to solve the unnaturalness problem in image restoration. More analysis is provided in the supplementary materials.
Figure 3: Overview structure of our method. For any existing restoration network, we train it with the aid of our DiffLoss to achieve higher performance and more natural visual results. The parameter of DiffLoss is frozen during training stage. During inference stage, we only have the optimized restoration network, without involving the DiffLoss.
@@ 4.1 Denoising Diffusion Probabilistic Models
Figure 4: Detailed design of DiffLoss. We devise our DiffLoss with $t$ step forward process and one step reverse process. The $t$ step forward process can be directly achieved with Eq. 2. After projecting the restored image and the clear image into intermediate noisy images with Eq. 2, these noisy images are further fed into the denoising UNet. The DiffLoss is designed to pull the output of the denoising UNet and the bottleneck feature closer.
@@ 4.3 Detailed design of DiffLoss
To this end, we redesign diffusion model delicately. We employ $t$ step forward diffusion process and one step reverse denoise step to minimize time-consuming as well as get symmetric image-image input-output pairs. Specifically, as shown in Fig. 3 and Fig. 4, we integrate the forward diffusion process and reverse denoise process. In the forward diffusion process, we get the intermediate noisy image $x_{t}$ via $q(x_{t}|x_{0},\varepsilon)$, which is expressed as Equation 7.
@@
Then comes the reverse denoise process. Note that Eq. 8 provides the way of reconstructing $x_{0}$ back. However, we need to implement Eq. 8 in a learnable way. Accordingly, we devise two ways by respectively replacing $x_{t}$ and $\varepsilon$ with the diffusion model learned ones, as shown in Fig. 3.
In the first way, starting form $x_{t}$, we get the pseudo Gaussian noise $\hat{\varepsilon}=f_{\theta}(x_{t},t)$. Then, the pseudo Gaussian noise map $\hat{\varepsilon}$ is employed to replace $\varepsilon$ in Eq. 8. The first inverse way $q(\hat{x_{0}}|x_{t},\hat{\varepsilon})$ is expressed as follows:
$\hat{x}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}x_{t}-\sqrt{(\frac{1}{\bar{\alpha}_{t}}-1)}\hat{\varepsilon}.$ (9)
The second manner starts from $x_{t}$ to $\hat{x}_{t-1}$ via $q(\hat{x}_{t-1}|x_{t},f_{\theta}(x_{t},t))$, and finally get the output $\hat{x_{0}}$ from $\hat{x}_{t-1}$ with Eq. 10:
$\hat{\hat{x}}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t-1}}}\hat{x}_{t-1}-\sqrt{(\frac{1}{\bar{\alpha}_{t-1}}-1)}\varepsilon,$ (10)
Figure 5: Visualization of the intermediate and output images of DiffLoss at different step t. Despite input a clear image into DiffLoss module, the intermediate stage reconstructed $\hat{x}_{0}$ and $\hat{\hat{x}}_{0}$ are not complete reconstructions of the input image.
Besides, we show $x_{t}$, $\hat{x}_{t-1}$, $\hat{x}_{0}$, and $\hat{\hat{x}}_{0}$ in Fig. 5 to visualize how the intermediate sampling stage changes with $t$. Note that although the visual difference between $x_{t}$ and $\hat{x}_{t}$ is negligible, the reconstructed $\hat{x}_{0}$ and $\hat{\hat{x}}_{0}$ are obviously different, and are not complete reconstruction of $x$. We can also direct constrain on $\hat{\varepsilon}$ and $\hat{x}_{t-1}$ without reconstructing back to $\hat{x}_{0}$ and $\hat{\hat{x}}_{0}$, which are work equally. In this work, we choose current way to better visually illustrate the difference of the intermediate sampling stages. In this way, the restoration networks can harness the generative capabilities of the diffusion model, as well as the sampling space’s reflection of the natural attributes of images, to obtain restored results with more naturalness.
If we decompose $f_{\theta}(x_{t},t)$ into an encoder $\mathcal{E}_{\theta}(\cdot)$ and a decoder $\mathcal{D}_{\theta}(\cdot)$ , then after obtaining the noisy version of clear image $\hat{x}_{t}$ and restored image $\hat{z}_{t}$, the form of h-space feature of them can be written as the following equations respectively:
$\hat{h}_{clr}^{t}=\mathcal{E}_{\theta}(x_{t},t),$ (11)
$\hat{h}_{rst}^{t}=\mathcal{E}_{\theta}(z_{t},t),$ (12)
where $\hat{h}_{clr}^{t}$ and $\hat{h}_{rst}^{t}$ reflect the semantic information of the clear image and restored image, respectively. By reducing the gap between these two terms, we can preserve more semantic information in the restored image.
@@
$L_{DiffLoss}$ $=\left\|\hat{x}_{0}-\hat{z}_{0}\right\|_{2}+\left\|\hat{\hat{x}}_{0}-\hat{\hat{z}}_{0}\right\|_{2}+\lambda\left\|\hat{h}_{clr}^{t}-\hat{h}_{rst}^{t}\right\|_{2},$ (13)
@@
where $\lambda$ and $\gamma$ are weight factors. We empirically set the weight $\lambda=0.01$ and $\gamma=0.001$. Besides, we also try adaptive weight which is correlated with timestep $t$. Both strategies perform similarly. A detailed ablation on the weight of DiffLoss is conducted in the experimental part. By employing these two strategies, the DiffLoss has the potential to enhance existing restoration methods, yielding restored images that are not only more natural but also reserve more semantic information.

### v2

# DiffLoss: unleashing diffusion model as constraint for training image restoration network

@@ Abstract
Image restoration aims to enhance low quality images, producing high quality images that exhibit natural visual characteristics and fine semantic attributes. Recently, the diffusion model has emerged as a powerful technique for image generation, and it has been explicitly employed as a backbone in image restoration tasks, yielding excellent results. However, it suffers from the drawbacks of slow inference speed and large model parameters due to its intrinsic characteristics. In this paper, we introduce a new perspective that implicitly leverages the diffusion model to assist the training of image restoration network, called DiffLoss, which drives the restoration results to be optimized for naturalness and semantic-aware visual effect. To achieve this, we utilize the mode coverage capability of the diffusion model to approximate the distribution of natural images and explore its ability to capture image semantic attributes. On the one hand, we extract intermediate noise to leverage its modeling capability of the distribution of natural images, which serves as a naturalness-oriented optimization space. On the other hand, we utilize the bottleneck features of diffusion model to harness its semantic attributes serving as a constraint on semantic level. By combining these two designs, the overall loss function is able to improve the perceptual quality of image restoration, resulting in visually pleasing and semantically enhanced outcomes. To validate the effectiveness of our method, we conduct experiments on various common image restoration tasks and benchmarks. Extensive experimental results demonstrate that our approach enhances the visual quality and semantic perception of the restoration network.
@@ 1 Introduction
In complex imaging environments, the quality of imaging often suffers from unpredictable degradations, such as low-light conditions, heterogeneous media, leading to information loss both in content and color and unnatural visual effects. Besides, these degradations also causes adverse effects on high-level computer vision tasks such as object detection [8, 35] and scene understanding [53, 52]. Image restoration aims to solve these issues by recovering high quality images from degradations to achieve better natural visual quality and fine semantic attributes. This task has been a longstanding and challenging problem, drawing the attention of researchers. Previous traditional methods utilized the inherent and general statistical property of degraded images to design algorithms with prior knowledge for image restoration[4, 24, 16, 15]. However, these hand-crafted image priors are drawn from specific observations with limited robustness, which may not be reliable for modeling the intrinsic characteristics of degraded images. With the development of deep learning, image restoration has made marvelous progress, where most of them rely on designing advanced architectures to learn degradation to clean mappings [34, 39, 14, 64, 21, 69]. However, due to the severe degradation and limited model capacity, the naturalness of the restored results is restricted by color and texture distortions, as shown in Fig. 1. In addition to further improving image visual quality, preserving semantic attributes is also significant for image restoration. Existing methods try to solve the problem by either involving multi-stage training or utilizing complex network structures[30, 67, 68], which result in inconvenience during implementation. With the emergence of the diffusion model, it has shown remarkable performance as a backbone in image restoration tasks with naturalness and semantic-aware visual results[10, 51, 63, 40, 29, 58]. However, these methods attempt to integrate diffusion models into the network architecture that has the limit of slow inference speed and accounting huge memory cost when deploying.
@@
Figure 1: (a): Illustration of the effect of our method to convert pixel constraint into distribution constraint.(b): Visual comparison between baselines with and without our DiffLoss. The top is produced with IAT[11] on LOL dataset[61] and the bottom is produced with MSBDN[14] on NH-HAZE dataset [2]. Previous loss is limited to the pixel level and suffers from unnaturalness problem, with color shift and content artifacts. Our DiffLoss leverages the powerful modeling capability of the diffusion model on the distribution of natural images, resulting in more natural outcomes in the image restoration process. Need to mention that DiffLoss is an optimization strategy. The improvement should be compared with the baseline method, instead of other restoration methods or the ground-truth images.
In this study, to address these issues, instead of creating new network architectures, we empower existing network frameworks with powerful prior on natural image distribution modeling and high-level semantic space. Inspiringly, the diffusion model has exhibited strong natural image generation capability [13] and possible semantic potential [3], which motivate us to exploit the "implicit" usage of the diffusion model and leverage it as an optimization prior to improve naturalness and semantic attributes for image restoration.
[…]
(b) Histogram of high-level feature distance of results on low-light image enhancement with synthesized data from ImageNet dataset.
[…]
In this paper, we propose a new perspective to resort these issues by introducing a naturalness-oriented and semantic-aware optimization mechanism by using diffusion model, dubbed DiffLoss, which has two parts. (1) Natural image distribution prior. Drawing inspiration from the diffusion model’s remarkable capacity to cover distributions in natural image generation, we leverage the Markov chain sampling characteristic of the diffusion model to project the restored results of existing networks into the noise sampling space and utilize it as a constraint to enhance the natural visual representation of images. Its improvement for natural visual performance can be observed in Fig. 1. (2) High-level semantic space prior. The bottleneck feature of diffusion models, also dubbed h-space feature, is verified to be natural high-level semantic space in [77, 23, 27] and is shown in Fig. 2. For both clear images and images with different degradations, we first add noise to them and then input them into the diffusion model. During the generation, we perform SVD decomposition on the feature in h-space and perturb its principal component. As the perturbation changes, the semantic appearance in the generated images also change. We extract the high-level features of these images in a ResNet50 that pre-trained on ImageNet for classification and measure the L2 distance of the perturbed image features from original image features. We find the perturbation in h-space can disrupt high-level features of clean images and restored images with our method. While for degraded images and restored images without our method, the disturbance is slightly, which implies degradations corrupt the semantic attributes of the images and DiffLoss can act as a constraint for semantic recovery in image restoration tasks.
@@
For implementation, we employ the unconditional diffusion model pre-trained on ImageNet dataset as optimization prior to exploit its immense clean data distribution. The diffusion model takes the restored image and ground truth as inputs and projects them into the distribution sampling space and semantic space by extract their sampling noise distribution and bottleneck feature. Within these spaces, the empowered restoration network can get more natural result and preserve more semantic attributes by pulling the intermediate sampling noise distribution and bottleneck feature of the restored image closer to that of the corresponding clear image. Different from previous methods that focus on dedicated model design or directly employ diffusion model as the restoration model, our DiffLoss works as a general and novel auxiliary training mechanism, which can endow existing restoration methods with both more natural and semantic-aware results with this effective training strategy. Additionally, it’s especially beneficial for empowering parameter-limited models as it involves naturalness prior. We also compared it with other loss functions that assist restoration networks, demonstrating excellent performance. Our method also involves no additional computations in the inference stage and easy for implementation. We verify the effectiveness of our method on substantial common image restoration tasks, including image dehazing, image deraining, and low-light image enhancement.
@@
  • We introduce the naturalness and semantic-aware modeling paradigm into restoration network by embedding the diffusion model as an auxiliary training mechanism, which has not been explored before. This training strategy alleviates the low quality issue caused by unnatural visual quality and lack of semantic attributes in existing methods.
  • Specifically, our approach leverages the fixed diffusion model to enable the extraction of intermediate sampling noise and semantic information, and it yields more natural and semantic-aware restoration when the DiffLoss is minimized.
  • Extensive experiments demonstrate that our DiffLoss empowers existing restoration methods compared with other loss functions, helps training efficient models and improves the classification ability on data with varying of degradation without involving additional computations in the inference stage at all.
@@ 2 Related work
2.1 Visual and semantic improvement for image restoration
Image restoration refers to the process of improving the quality of a degraded or damaged image by removing various types of degradations. With the advent of deep learning, numerous methods leveraging deep neural networks for image restoration have emerged, such as low-light enhancement[60, 73, 74, 78], super-resolution[33, 38, 75, 76], dehazing[4, 7, 15, 24, 36, 34, 71], deraining[19, 37, 41, 45, 48], deblurring[5, 9, 20, 32, 49]. However, the severe degradation limits the naturalness of restored results, causing color and texture distortions. Moreover, it is proved that solely relying on visual quality metrics during the restoration process without considering the semantic aspects of quality will negatively impacted performance in downstream tasks. Some efforts [30, 67, 68] have attempted to improve the semantic attributes of restored network output images through time-consuming training strategies and complex networks, but they often result in inconvenience and redundancy in practical applications.
@@ 2.2 Diffusion models for image restoration
However, these methods aim to incorporate diffusion models as backbone, which suffer from the drawbacks of slow inference speed and significant memory consumption during deployment. We circumvent these limitations from a new perspective by exploring diffusion model as an auxiliary training mechanism to empower the learning capability of existing image restoration networks, without involving additional computations in the inference stage.
Figure 3: Overview structure of our method. The parameter of DiffLoss is frozen during training stage. For any existing restoration network, we train it with the aid of our DiffLoss to achieve higher natural visual and semantic performance. More implementation details of DiffLoss can be found in Fig. 4. During inference stage, we only have the optimized restoration network, without involving the DiffLoss.
@@ 3.1 Denoising Diffusion Probabilistic Models
Figure 4: Detailed design of DiffLoss. We devise our DiffLoss with $t$ step forward process and one step reverse process. The $t$ step forward process can be directly achieved with Eq. 7. Then we projecting these noisy images into intermediate noise with Eq. 4 after fed into the denoising UNet. We also get h-space vector from bottleneck of the UNet,which contains semantic information, as described in Eq. 12 and Eq. 13. The DiffLoss is designed to pull the output of the denoising UNet and the bottleneck feature closer.
@@ 3.3 Detailed design of DiffLoss
To this end, we redesign diffusion model delicately. We employ $t$ step forward diffusion process and one step reverse denoise step to minimize time-consuming as well as get symmetric image-image input-output pairs. Specifically, as shown in Fig. 3 and Fig. 4, we integrate the forward diffusion process and reverse denoise process. In the forward diffusion process, we get the intermediate noisy image $x_{t}$ via $q(x_{t}|x_{0},\varepsilon)$. Note that $t$ is obtained through uniform distribution sampling, which is expressed as Eq. 7. And Eq. 8 provides the way of reconstructing $x_{0}$ back:
@@
Then comes the reverse denoise process. We need to implement Eq. 8 in a learnable way. Accordingly, we devise a way by respectively replacing $\varepsilon$ with the diffusion model learned ones, as shown in Fig. 3.
Starting from clear image with added noise $x_{t}$, we get the pseudo Gaussian noise $\hat{\varepsilon}^{clr}_{t}=f_{\theta}(x_{t},t)$. Then, the pseudo Gaussian noise map $\hat{\varepsilon}^{clr}_{t}$ is employed to replace $\varepsilon$ in Eq. 8. The inverse way $q(\hat{x_{0}}|x_{t},\hat{\varepsilon}^{clr}_{t})$ is expressed as follows:
$\hat{x}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}x_{t}-\sqrt{(\frac{1}{\bar{\alpha}_{t}}-1)}\hat{\varepsilon}^{clr}_{t}.$ (9)
Similarly, the output results of the restoration network can undergo the same process to obtain the following equation.
$\hat{z}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}z_{t}-\sqrt{(\frac{1}{\bar{\alpha}_{t}}-1)}\hat{\varepsilon}^{rst}_{t}.$ (10)
Note that we can also obtain $\hat{x}_{t-1}$ and $\hat{z}_{t-1}$ using Eq. 4. We choose to direct constrain on $\hat{\varepsilon}^{clr}_{t}$ and $\hat{\varepsilon}^{rst}_{t}$ without reconstructing back to $\hat{x}_{0}$ and $\hat{z}_{0}$ or using $\hat{x}_{t-1}$ and $\hat{z}_{t-1}$ as constraint, cause it has best performance and is shown in Sec. 4.3. The loss function can be expressed as follows:
$L_{nat}$ $=\left\|\hat{\varepsilon}^{clr}_{t}-\hat{\varepsilon}^{rst}_{t}\right\|_{2}.$ (11)
In this way, the restoration networks can harness the generative capabilities of the diffusion model, as well as the sampling space’s reflection of the natural attributes of images, to obtain restored results with more naturalness.
If we decompose $f_{\theta}(x_{t},t)$ into an encoder $\mathcal{E}_{\theta}(\cdot)$ and a decoder $\mathcal{D}_{\theta}(\cdot)$ , then after obtaining the noisy version of clear image $\hat{x}_{t}$ and restored image $\hat{z}_{t}$, the form of semantic feature of them can be written as the following equations respectively:
$\hat{h}_{t}^{clr}=\mathcal{E}_{\theta}(x_{t},t),$ (12)
$\hat{h}_{t}^{rst}=\mathcal{E}_{\theta}(z_{t},t),$ (13)
where $\hat{h}_{t}^{clr}$ and $\hat{h}_{t}^{rst}$ reflect bottleneck features in the middle layers of U-Net from the clear image and restored image, respectively. By reducing the gap between these two terms, we can preserve more semantic information in the restored image, which is expressed as follows:
$L_{sem}$ $=\left\|\hat{h}_{t}^{clr}-\hat{h}_{t}^{rst}\right\|_{2}.$ (14)
@@
$L_{DiffLoss}$ $=L_{nat}+\lambda L_{sem},$ (15)
@@
where $\lambda$ and $\gamma$ are weight factors. We set the weight $\lambda=0.01$ and $\gamma=0.001$, which is discussed in Sec. 4.3. Besides, we also try adaptive weight which is correlated with timestep $t$. Both strategies perform similarly. By employing these two strategies, the DiffLoss has the potential to enhance existing restoration methods, yielding restored images that are not only more natural but also reserve more semantic information.